OpenStudy (anonymous):

Find two consecutive odd integers such that 1/3 the smaller plus twice the larger equals 7 more than the sum of the two numbers. @Algebra I

OpenStudy (anonymous):

See like i know its starts like: x, x+2 x+x+2=1/3x + 2(x+2)........... or something of that nature :P

OpenStudy (anonymous):

This is a system of equations. x and x+2 x/3+2(x+2)+7=x+x+2 x/3+2x+4+7=2x+2 7x/3+11=2x+2 x/3=-9 x=-27 x+2=-25